OpenStudy (anonymous):

Given a soda can with a volume of 36 and a diameter of 4, what is the volume of a cone that fits perfectly inside the soda can?

OpenStudy (cwrw238):

volume of the can = pi r^2h = 36 r = 1/2 * 4 = 2 so 36 = pi*2^2 * h h = 36 / 4 pi

OpenStudy (cwrw238):

given that volume of the cone thet fits into the can = (1/3) pi r^2 h can you find this volume? h = 36/4pi = 9/pi and r = 2

OpenStudy (anonymous):

2.9

OpenStudy (cwrw238):

oh - there was no need for me to do all this!!

OpenStudy (anonymous):

9/pi is 2.9 when rounded to the nearest tenth

OpenStudy (anonymous):

and thank you so much!

OpenStudy (cwrw238):

as the volume of the cone = 1/3 * volume of the cylinder the answer is 1/3 * 36 = 12

OpenStudy (anonymous):

so the final answer would be 12?

OpenStudy (cwrw238):

yes - there was no need to do all those calculations!! - the volume of the cone is 1/3 times the volume of the cylinder

OpenStudy (cwrw238):

V can = pi r^2 h V cone = (1/3 pi r^2 h r and h is the same for both can and cone

OpenStudy (cwrw238):

I don't know why I didn't see that in the first place!!

OpenStudy (cwrw238):

do you follow what wrote?

OpenStudy (cwrw238):

its not 2.9 9/pi is the height of the the can ( and the cone)

OpenStudy (anonymous):

okay, the final answer is 12

OpenStudy (anonymous):

I think. Haha

OpenStudy (cwrw238):

if we substitute h = 9/pi into the formula for the volume of the cone we get 12 which confirms the result

OpenStudy (anonymous):

Okay. Thank you so much! (:

OpenStudy (cwrw238):